OpenStudy (abbycross167):

Use your equation from #3 and show your work to predict the European size of Shaq O’Neal’s U.S. size of 22. Here is the equation from question #3~ y=5/2x+26.5

OpenStudy (abbycross167):

@Directrix can you please help me?

Directrix (directrix):

Where is the original problem posted? I need to see what x and y represent and why the equation has this in it: ~ abbycross167

Directrix (directrix):

@abbycross167

OpenStudy (abbycross167):

here is the graph with the sizes~ U.S size 1 3 5 7 9 11 European size 31 34 36 39 41 44

OpenStudy (abbycross167):

Equation = y=5/2x+26.5 Plug x=8 into the equation Y=5/2(8)+26.5 Y= 46.5 m = (y2 – y1) / (x2 – x1) m = (39-34) / (5-3) m = 5/2 b=26.5 Equation = y=5/2x+26.5

OpenStudy (abbycross167):

@Directrix was that what you wanted?

Directrix (directrix):

Yes. Thanks. I will check the equation before I do the size conversion.

Directrix (directrix):

1 3 5 7 9 11 31 34 36 39 41 44 To get the equation, you used the points (5,39) and (3,34). They do not match the chart. 5 goes with 36 and (3,34) Write the equation again using (5,36) and (3,34). Read the chart of values carefully.

Directrix (directrix):

I used points (1,31) and (3,34) to get the equation as: y = (3/2) x + 59/2 y = (3/2) * 22 + 59/2 y = 62.5 for Shaq's European shoe size Now, I am worried if numbers in the table are out of order or incorrect. Maybe you could post a screenshot. Do check the posted points.I If the chart is off kilter, then the calculations will be wrong. @abbycross167

OpenStudy (abbycross167):

@Directrix could you show me how you got the equation y = (3/2) * 22 + 59/2? Like show me the steps? please?

OpenStudy (abbycross167):

Thank you very much btw!!!

Directrix (directrix):

The slope of the line, m, is found from two points from the chart: (1,31) and (3,34). m = (34-31)/(3-1) = 3/2 With slope 3/2 and using selected point (1,31) and the slope-point form of a line, y - 31 = (3/2)*(x - 1) y - 31 = (3/2)*x - 3/2 y = (3/2)*x - 3/2 + 31 y = (3/2)*x -3/2 + 62/2 y = (3/2)*x + 59/2

Directrix (directrix):

Shaq wears shoe size US of 22. Replace x by 22 in y = (3/2)*x + 59/2 y = (3/2) * 22 + 59/2 y = 62.5 for Shaq's European shoe size

OpenStudy (abbycross167):

Ok Awsome!! Thank you so much!!

Directrix (directrix):

You are welcome.